OpenStudy (anonymous):

One number is four times another number. Their sum is 20. What are the two numbers a.5 and 20 b.4 and 16 c.5 and 15 d.2 and 18

OpenStudy (anonymous):

i got b never mind:)

mathslover (mathslover):

ok ! let the number be x : the other number will be 4x given x + 4x = 20 5x = 20 \(\huge{x=\frac{20}{5}=4}\) hence x = 4 , 4x = 4 * 4 = 16

mathslover (mathslover):

u r right congrats :)